OpenStudy (anonymous):

finding the slope of the tangent with e^9 in it??

OpenStudy (anonymous):

i got 8e^9..but how would i write that in exact decimal form?

OpenStudy (nincompoop):

you don't have to

OpenStudy (anonymous):

i put 8e^(9)x , and also e^(9)x.. but it says its wrong! :(

zepdrix (zepdrix):

Hmm I came up with something very different for the slope at e^9..

zepdrix (zepdrix):

You're doing implicit differentiation?

OpenStudy (anonymous):

i just found the derivative andpluged in (1, e^9) for x and y

OpenStudy (nincompoop):

first, just get the general tangent formula for the equation you were given. it's implicitly done then evaluate it at the point you were given

zepdrix (zepdrix):

Did you remember to apply the `product rule` to the middle term?

zepdrix (zepdrix):

Yes, that's a good approach skull, let's make sure your derivative is correct though :)

OpenStudy (anonymous):

i got ((9x-1)y)/x for the derivative

OpenStudy (anonymous):

\[(9x-1)y/1\]

OpenStudy (nincompoop):

can you show us all the steps?

OpenStudy (anonymous):

ln(xy) = 9x e^[ln(xy)] = e^(9x) xy = e^(9x) y = [e^(9x)] / x y' = [e^(9x) * (9x - 1)] / (x^2) y'(x = 1) = [e^(9*1) * (9*1 - 1)] / (1^2) = 8*e^9

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

nvm it was asking for the equation! :)